OpenStudy (anonymous):

x+y=2 and 2x+5y=16 what is the solution

OpenStudy (anonymous):

Do you just want the answer or the work to solve it

OpenStudy (anonymous):

work so i could solve other ones

OpenStudy (anonymous):

x+y=2 and 2x+5y=16 for the variables x and y. This means, see if it can be simplified at all before attempting to solve it. x+y evaluates to x+y So, all-in-all, your first equation can be written as: x+y = 2 ---------------------------------------------------------------- Now, let's work on your second equation,2x+5y=16 Multiply x and 2 Multiply x and 1 answer is x 2*x evaluates to 2x Multiply y and 5 Multiply y and 1 The y just gets copied along. The answer is y roblem to solve is: The goal here is to solve x+y=2 and 2x+5y=16 for the variables x and y. First, let's work on your first equation,x+y=2 This means, see if it can be simplified at all before attempting to solve it. x+y evaluates to x+y So, all-in-all, your first equation can be written as: x+y = 2 Now, let's work on your second equation,2x+5y=16 Multiply x and 2 Multiply x and 1 The x just gets copied along. The answer is x x 2*x evaluates to 2x Multiply y and 5 Multiply y and 1 (The y just gets copied along.) The answer is y y 5*y evaluates to 5y 2*x+5*y evaluates to 2x+5y So, all-in-all, your second equation can be written as: 2x+5y = 16 After this initial survey of the equations, the system of equations we'll set out to solve is: x+y = 2 and 2x+5y = 16 Let's start by solving x+y = 2 for the variable x. Move the y to the right hand side by subtracting y from both sides, like this: From the left hand side: y - y = 0 The answer is x From the right hand side: The answer is 2-y nowwww u solve 2x+5y = 16 for the variable y. Move the 2x to the right hand side by subtracting 2x from both sides From the left hand side: 2x - 2x = 0 The answer is 5y From the right hand side: The answer is 16-2x now its 5y = 16-2x The last step is to divide both sides of the equation by 5 like this: To divide y by 1 The y just gets copied along in the numerator. The answer is y 5y ÷ 5 = y To divide 16-2x by 5 divide each term in 16-2x by 5 term by term. 16 ÷ 5 = 16/5 To divide x by 1 The x just gets copied along in the numerator. The answer is x -2x ÷ 5 = -2/5x The solution to your equation is: y =16/5-2/5x Now, plug the earlier result, x=2-y, in for x everywhere it occurs in y=16/5-2/5x. his gives y=16/5-2/5(2-y). Now all we have to do is solve this for y,to have our first solution. 16/5 evaluates to 16/5 2/5 evaluates to 2/5 2-y evaluates to 2-y Multiply 2/5 by 2-y we multiply 2/5 by each term in 2-y term by term. Multiply 2/5 and 2 1 2/5 × 2 =4/5 Multiply 2/5 and -y Multiply 1 and y The y just gets copied along. Yeah you get the point just keep bringing it down then Lastly, to find the solution for x, we plug this answer for y into the earlier result that x=2-y. This gives x=2-(4). Now, simplify this. 2-(4) evaluates to -2 x= -2 So, the solutions to your equations are: x= -2 and y= 4

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Anytime !